OpenStudy (anonymous):

The expression | 2 + 4n | - 4 = 14 produces the answers n = 4 and n = 5. TRUE OR FALSE

OpenStudy (anonymous):

true :)

Directrix (directrix):

I got -5 and 4. So, FALSE but let me check my work.

OpenStudy (anonymous):

oh wait that's what I got too... sorry I thought it was negative 5... I didn't see that it was positive 5.... directrix is right :)

Directrix (directrix):

@ madscientist --> I made those same type sign-us errors twice earlier today. I had to take a break and put drops in my eyes. No worries.

Directrix (directrix):

Where's ChinaGurl?

OpenStudy (anonymous):

| 2 + 4n | - 4 = 14 | 2 + 4n | = 14 + 4 = 18 2 + 4n = 18 4n = 18 -2 = 16 ->n = 16/4 = 4 2+ n = -18 -> n = -18 + 2 = -16

OpenStudy (anonymous):

Oop 2 + 4n = -18 4n = -18 - 2 = -20 -> n = -20/4 = -5 Thus n = 4, n = -5

OpenStudy (anonymous):

Yes, my screen keeps fluctuate, it causes typo many times!